OpenStudy (anonymous):

Is the function below exponential? G(t)= (4 . t)^4 If so, write the function in the form G(t)=ab^t and find the value for a and b. How do I do this question?

zepdrix (zepdrix):

Is this what the function looks like?\[\Large\rm G(t)=(4t)^4\]

OpenStudy (anonymous):

G(t) = (4*t)^4

zepdrix (zepdrix):

Multiplication..? Yah so that's implied in what I wrote..

OpenStudy (anonymous):

well it shows up as a dot. So I imagine it is a multiplication

zepdrix (zepdrix):

So we need to see whether or not we can get it into this form:\[\Large\rm G(t)=\color{orangered}{a}\color{royalblue}{b}^t\]

zepdrix (zepdrix):

Remember your exponent rule?\[\Large\rm (ab)^c=a^c~b^c\]

zepdrix (zepdrix):

Applying this rule gives us:\[\Large\rm G(t)=4^4~t^4\]

zepdrix (zepdrix):

Hmm looks like this is a `polynomial function`, not an exponential one. See how our variable is the `base`? Unless we're misinterpreting the dot lol

OpenStudy (anonymous):

Okay so to be exponential we would need the variable in the c position (not the base)

zepdrix (zepdrix):

ya c:

OpenStudy (anonymous):

And since it is not, then it is a polynomial function.

OpenStudy (anonymous):

Great thanks again.

zepdrix (zepdrix):

Since it's not, that implies it is "something else". But yes, it turned out to be a polynomial in this case :) Since the exponent was a positive integer.